When the price of a good x falls from £5 to £3, the demand for good y increases from 14 units to 18 units. Use the mid-point approach to calculate the cross price elasticity of demand and identify whether the goods are complements or substitutes. (4 marks)


1
Expert's answer
2021-01-13T11:48:08-0500

"\%" change in quantity "=\frac{Q_{2}-Q_{1}}{(Q_{2}+Q_{1})\div 2}\times 100\%=\frac{(18-14)}{(18+14)\div 2}\times 100=4/16\times100=25\%"

"\%" change in price "=\frac{(P_{2}-P_{2})}{(P_{2}+P_{1})\div 2}\times 100=\frac{(3-5)}{(3+5)\div 2}\times 100=-2/4\times100=-50\%"

cross price elasticity of demand"=\%" change in quantity"/\%" change in price"=25\/-50=-0.5"

Because the cross-price elasticity is negative, good x and good y are complementary goods.